Consider the following chemical equation 2Na3PO4(aq) + 3CuCl2(aq) -> Cu3(PO4)2(s)+6NaCl (aq)

a. What volume in mL of .1950M Na3PO4 is necessary to completly react with 85,6mL of .205 M CuCl2?

b. If 2.100 g are actually produced in an experiment what is the % reaction yield?
Answer #1

a) .205 M = x mol of CuCl2 / .0856L

    .01755 mol of CuCl2 = 3

    x mol of Na3PO4     =     2

.0117 mol of Na3PO4 / L = .195M

V = 60 mL of Na3PO4

b) If we take as a theorical yield 1 mol, the yield is = 2.1 g / 380.58 g = .55%

If we take the previous experiment: 2.1g / 2.2264 g = 94.32%